OpenStudy (anonymous):

integral (e^(9x))/(e^(18x)-e^(9x))dx

hartnn (hartnn):

did you try to put u = e^(9x) ? du = ... ?

OpenStudy (anonymous):

i tried using substitution and then used partial fractions, but my answer was still wrong

OpenStudy (dumbcow):

\[\rightarrow \frac{1}{9} \int\limits \frac{du}{u^2 -u} = \frac{1}{9} \int\limits \frac{du}{u(u-1)} = \frac{1}{9} \int\limits \frac{du}{u -1} - \frac{1}{9} \int\limits \frac{du}{u}\] \[= \frac{1}{9}\ln (u-1) -\frac{1}{9} \ln u \] \[=\frac{1}{9} \ln (e^{9x} -1) - x\]

OpenStudy (anonymous):

i got the same answer, but the system read it as incorrect, so i think there might be something wrong with the system itself, it happens very often.

hartnn (hartnn):

maybe you forgot to add the "+c" in the end :P

OpenStudy (anonymous):

i tried to it with +constant and without so. i'll just email my teacher and confirm that i have my answer right.and get the point for that problem

hartnn (hartnn):

good luck!